# Adaptive resolution of fine scales in modes of microstructured optical fibers

### v1

@@ 3. The error estimator for eigenvalues
with $a,b,c,d,m$ as in (13). Viewing $u$ as a right eigenfunction, we also assume that there is a corresponding left eigenfunction $0\neq\tilde{u}\in H$ satisfying
@@
(20b) $\rho_{T,2}^{2}$ $=h_{T}^{2}\left\|\operatorname{div}(\gamma\operatorname{grad}\varphi_{h}+V\gamma E_{\tau,h}-Z_{h}^{2}\gamma E_{h})\right\|_{T}^{2}$
$\quad+\frac{h_{T}}{2}\left\|\left\llbracket(\gamma\operatorname{grad}\varphi_{h}+V\gamma E_{\tau,h}-Z_{h}^{2}\gamma E_{h})\cdot\nu\right\rrbracket\right\|_{\partial T\setminus\partial\Omega}^{2},$
@@
The adaptive algorithm we use is motivated by the next result (Theorem 1 below), which gives a global reliability estimate for the error indicator $\eta_{T}.$ We prove it using the DWR approach laid out in [1, 12, 26], but bring in an additional ingredient to decompose $H(\operatorname{curl})$ approximation errors into locally controllable gradients and a regular remainder using the results of [27]. As usual, we make a “saturation assumption,” which here takes the form that the discretizations considered are fine enough so that
(21) $\left(\max_{x\in\Omega}\|\gamma(x)\|_{2}\right)\|E_{\tau}-E_{\tau,h}\|_{H(\operatorname{curl},\Omega)}\|\tilde{E}_{\tau}-\tilde{E}_{\tau,h}\|_{H(\operatorname{curl},\Omega)}<\frac{1}{2}$
@@
Theorem 1.
Suppose $u=(E_{\tau},\varphi)$, $u_{h}=(E_{\tau,h},\varphi_{h}),$ $\tilde{u}=(\tilde{E}_{\tau},\tilde{\varphi})$, $\tilde{u}_{h}=(\tilde{E}_{\tau,h},\tilde{\varphi}_{h})$, solve (16), (17), (18) and (19), respectively, with accompanying exact and discrete eigenvalues $\lambda$ and $\lambda_{h}$. Suppose also that (21) holds. Then the error in eigenvalue can be bounded by the error indicators defined in (20) by
@@
$+A(u,\tilde{u}_{h})-\lambda_{h}B(u,\tilde{u}_{h}),$ by (18) and (19),
$=(\lambda-\lambda_{h})\big{[}B(u_{h},\tilde{u})+B(u,\tilde{u}_{h})\big{]},$ by (16) and (17),
$=(\lambda-\lambda_{h})\big{[}B(u_{h}-u,\tilde{u}-\tilde{u}_{h})+2\big{]},$ by (17b) and (19b),
@@
$\qquad\qquad+(\varphi-I_{h}\varphi,(\bar{\gamma}\tilde{E}_{\tau,h})\cdot\nu)_{\partial T\setminus\partial\Omega}-(\varphi-I_{h}\varphi,\operatorname{div}(\bar{\gamma}\tilde{E}_{\tau,h})+n_{2}^{2}\bar{\zeta}\bar{\eta}^{\prime}\tilde{\varphi}_{h})_{T}\Big{]},$
@@
$\lesssim\sum_{T\in\Omega_{h}}(h_{T}^{-1}\|z\|+\|\operatorname{grad}z\|_{T})\;\tilde{\rho}_{T,1}\;\lesssim\;\sum_{T\in\Omega_{h}}\|\operatorname{curl}E_{\tau}\|_{\Omega_{T}}\;\tilde{\rho}_{T,1},$
@@
$\tilde{\rho}(u-v_{h})\lesssim\sum_{T\in\Omega_{h}}\omega_{T,1}\;\tilde{\rho}_{T,1}+\omega_{T,2}\;\tilde{\rho}_{T,2}+\omega_{T,3}\;\tilde{\rho}_{T,3}.$
@@
$\rho(u-\tilde{v}_{h})\lesssim\sum_{T\in\Omega_{h}}\tilde{\omega}_{T,1}\;\rho_{T,1}+\tilde{\omega}_{T,2}\;\rho_{T,2}+\tilde{\omega}_{T,3}\;\rho_{T,3}.$
Using these estimates in (23) and noting that $21-\sigma>1/$ 2 due to the saturation assumption (21), the proof is completed. ∎
@@
This is used in Algorithm 1, which outlines the adaptive strategy we implemented.
@@
Input: Initial mesh $\Omega_{h}^{(0)},$ refinement threshold parameter $\theta$, an eigenvalue search region $D\subseteq\mathbb{C}$, and the maximum permitted number of degrees of freedom $N_{\text{max}}$.
@@
4       Assemble the sparse matrices for (18), namely $A_{h},B_{h}$, using $\Omega_{h}^{({\ell})}$ ;
5        // SOLVE
6
7      Use a sparse eigensolver to compute all eigenvalues $\lambda_{h}^{({\ell},k)}\equiv(Z_{h}^{({\ell},k)})^{2}$ of (18), within the search region $D$, corresponding eigenvectors $u_{h}^{({\ell},k)}$ satisfying
$A_{h}u_{h}^{({\ell},k)}=\lambda_{h}^{({\ell},k)}B_{h}u_{h}^{({\ell},k)},$
and the associated left eigenvectors $\tilde{u}_{h}^{({\ell},k)}$, for $k=1\ldots,K_{\ell}$. Collect these eigenvalues into a set $\varLambda_{h}^{({\ell})}=\{\lambda_{h}^{({\ell},k)}:\,k=1,\dots,K_{\ell}\}$;
8
9      for $k\leftarrow 1$ to $K_{\ell}$ do // ESTIMATE
10             Use (26) to compute the element-wise error indicators $\eta_{T}^{({\ell},k)}\equiv\eta_{T}^{(h)}(\lambda_{h}^{({\ell},k)},u_{h}^{({\ell},k)},\tilde{u}_{h}^{({\ell},k)})$ for each element $T\in\Omega_{h}^{({\ell})}$;
@@ 4       Assemble the sparse matrices for (18), namely $A_{h},B_{h}$, using $\Omega_{h}^{({\ell})}$ ;
12      for $T\in\Omega_{h}^{({\ell})}$ do
13             $\eta_{T}^{({\ell})}\leftarrow\displaystyle{\max_{k=1,\dots,K_{\ell}}}\eta_{T}^{({\ell},k)}$;
@@
15      for $T^{\prime}\in\Omega_{h}^{({\ell})}$ do // MARK
16             if $\eta_{T^{\prime}}^{({\ell})}>\textnormal{{$\theta$}}\ \displaystyle{\max_{T\in\Omega_{h}}}\,\eta_{T}^{({\ell})}$ then
17                   mark element $T^{\prime}$ for refinement;
18
19
20
21      Refine marked elements in $\Omega_{h}^{({\ell})}$ (as well as surrounding elements as needed to obtain conformity) to generate the next mesh $\Omega_{h}^{({\ell}+1)}$ ;
22        // REFINE
23
24      if number of degrees of freedom on $\Omega_{h}^{({\ell}+1)}>N_{\text{max}}$ then
25             break;
26
27
28return $\Omega_{h}^{({\ell})}$ and $\varLambda_{h}^{({\ell})}$;
@@
Algorithm 1 can use any sparse eigensolver. In our computations, we use the FEAST contour integral eigensolver [25, 9] (whose specific application to leaky modes is detailed in [10, Section A.4]), which takes as input a search region enclosed by a simple closed contour (denoted by $D$ in the algorithm). We use circular contours of small enough radius centered at $\hat{\lambda}_{h}^{(n)}=\frac{1}{K_{n}}\sum_{k=1}^{K_{n}}\lambda_{h}^{(n,k)}$, i.e., these centers track changes within each adaptive iteration and $D$ may therefore vary with iteration number $\ell$. How aggressively the refinements are made is determined by the parameter $\theta$. For our computations, we used $\theta=0.5$, a value determined by trial and error to give us a good balance between localization of the refinements and a steady increase in the number of degrees of freedom. So as to represent all eigenmodes of the eigenvalue cluster on the same mesh, the algorithm uses a marking strategy based on the maximum of the error indicators generated by each eigenmode. Other marking strategies are possible, as in [32, 1, 12]. As a stopping criterion, we set the maximum number of degrees of freedom to $N_{\text{max}}=2\times 10^{6}$ in our numerical experiments. This means that the algorithm will halt if the number of degrees associated with the $\ell$-th mesh exceeds $N_{\text{max}}$.
4. Verification by Bragg fiber
@@
These prior studies make the Bragg configuration an ideal example for verifying our adaptive strategy. Because we can semi-analytically compute the exact eigenvalue $\lambda=Z^{2}$ for this geometry, we are able to study the history of actual eigenvalue errors as the adaptive mesh refinement proceeds. Moreover, starting with an unbiased mesh that disregards the above-mentioned prior knowledge of the fine-scale features in the modes, we investigate if the adaptivity process is able to sense these features automatically and guide refinement to capture them.
@@
(b) Final mesh for the Bragg fiber, after twenty-four adaptive iterations.
@@
Figure 2. Results from the adaptive algorithm applied to a Bragg fiber.
In our numerical experiments for this fiber, we use the following settings for parameters (recalling that the parameters were defined in Section 3). The length scale in nondimensionalization is set to $L=1.5\times 10^{-5}$ m. The operating wavelength is determined by setting $k=\frac{2\pi}{1.7}\times 10^{6}\leavevmode\nobreak\ \text{m}^{-1}$. The refractive index is a piecewise-constant function whose values are $n_{\rm air}=1.00027717$ in the air regions and $n_{\rm glass}=1.43881648$ in the glass ring. The following non-dimensionalized radii are used: the air core has radius $r_{\rm core}=2.7183$; the glass ring starts at $r_{\rm core}$ and has an outer radius $r_{\rm outer}=3.385$, with a thickness of $t_{\rm ring}=0.66666667$; the outer air region starts at $r_{\rm outer}$ and extends to $r_{0}=4.385$, with a thickness of $t_{\rm air}=1.0$; and the PML starts at $r_{0}$ and extends to $r_{1}=8.05166666$, with a thickness of $t_{\rm PML}=3.66666667$. PML strength (see (14)) is set to $\alpha=2.0$. We set polynomial degree of $p=6$ to define the discretization (15).
@@
Next, we turn to examine the accuracy of the computed eigenvalue. The results are in Figure 3. The exact nondimensional eigenvalue we aim to approximate, obtained from semi-analytical computations, is $Z^{2}=0.80953881+0.00170153i$, and the exact eigenvalue cluster is the singleton containing it, $\varLambda=\{Z^{2}\}$. The discrete eigenvalue cluster $\varLambda_{h}$ is computed by providing a circular contour to FEAST centered around the exact eigenvalue. When the eigensolver converged, $\varLambda_{h}$ generally had two elements, matching the multiplicity of the exact eigenvalue. Let the Hausdorff distance between two sets $\Upsilon_{1}$ and $\Upsilon_{2}$ be denoted by $d(\Upsilon_{1},\Upsilon_{2})=\max(\sup_{\gamma_{1}\in\Upsilon_{1}}\operatorname{dist}(\gamma_{1},\Upsilon_{2}),\sup_{\gamma_{2}\in\Upsilon_{2}}\operatorname{dist}(\gamma_{2},\Upsilon_{1}))$. We report the convergence of the discrete eigenvalues in $\varLambda_{h}$ to $\varLambda$ using the Hausdorff distance in Figure 3.
@@
Figure 3. Convergence history of the cluster of discrete eigenvalues to the exact eigenvalue for the Bragg fiber, during the adaptive mesh refinement.
@@
$10^{5}$$10^{6}$$10^{-5}$$10^{-3}$$10^{-1}$$10^{1}$umber of degrees of freedom at the $\ell$-th iteration$|\text{CL}_{\ell}-\operatorname{CL}|$
Figure 4. Convergence history of the confinement loss for the Bragg fiber, during the adaptive mesh refinement. Here, $\operatorname{CL}(\beta)=\operatorname{Im}(\beta)\times 20/\ln(10)$ is the confinement loss in decibels per meter.
5. Three microstructured fibers

### v2

@@ 3 The error estimator for eigenvalues
with $a,b,c,d,m$ as in (13). Viewing $u$ as a right eigenfunction, we also assume that there is a corresponding left eigenfunction $0\neq\tilde{u}=(\tilde{E}_{\tau},\tilde{\varphi})\in H$ satisfying
@@
(20b) $\rho_{T,2}^{2}$ $=h_{T}^{2}\left\|\operatorname{div}(\gamma\operatorname{grad}\varphi_{h}+V\gamma E_{\tau,h}-Z_{h}^{2}\gamma E_{\tau,h})\right\|_{T}^{2}$
$\quad+\frac{h_{T}}{2}\left\|\left\llbracket(\gamma\operatorname{grad}\varphi_{h}+V\gamma E_{\tau,h}-Z_{h}^{2}\gamma E_{\tau,h})\cdot\nu\right\rrbracket\right\|_{\partial T\setminus\partial\Omega}^{2},$
@@
The adaptive algorithm we use is motivated by the next result (Theorem 3.1 below), which gives a global reliability estimate for the error indicator $\eta_{T}$. We prove it using the DWR approach laid out in [1, 12, 26], but bring in an additional ingredient to decompose $H(\operatorname{curl})$ approximation errors into locally controllable gradients and a regular remainder using the results of [27]. As usual, we make a “saturation assumption,” which here takes the form that the discretizations considered are fine enough so that
(21) ${\left(\max_{x\in\Omega}\|\gamma(x)\|_{2}\right)\,\|E_{\tau}-E_{\tau,h}\|_{\Omega}\;\|\tilde{E}_{\tau}-\tilde{E}_{\tau,h}\|_{\Omega}\,<1}$
@@
Theorem 3.1.
Suppose the vectors $u=(E_{\tau},\varphi)$, $u_{h}=(E_{\tau,h},\varphi_{h}),$ $\tilde{u}=(\tilde{E}_{\tau},\tilde{\varphi})$, $\tilde{u}_{h}=(\tilde{E}_{\tau,h},\tilde{\varphi}_{h})$, solve (16), (17), (18) and (19), respectively, with accompanying exact and discrete eigenvalues $\lambda$ and $\lambda_{h}$. Suppose also that (21) holds. Then the error in eigenvalue can be bounded by the error indicators defined in (20) by
@@
$+A(u,\tilde{u}_{h})-\lambda_{h}B(u,\tilde{u}_{h}),$ by (18) & (19),
$=(\lambda-\lambda_{h})\big{[}B(u_{h},\tilde{u})+B(u,\tilde{u}_{h})\big{]},$ by (16) & (17),
$=(\lambda-\lambda_{h})\big{[}B(u_{h}-u,\tilde{u}-\tilde{u}_{h})+2\big{]},$ by (17b) & (19b),
@@
$\qquad\qquad+(\varphi-I_{h}\varphi,(\bar{\gamma}\tilde{E}_{\tau,h})\cdot\nu)_{\partial T\setminus\partial\Omega}$
$\qquad\qquad-(\varphi-I_{h}\varphi,\operatorname{div}(\bar{\gamma}\tilde{E}_{\tau,h})+n_{2}^{2}\bar{\zeta}\bar{\eta}^{\prime}\tilde{\varphi}_{h})_{T}\Big{]},$
@@
$\lesssim\sum_{T\in\Omega_{h}}(h_{T}^{-1}\|z\|+\|\operatorname{grad}z\|_{T})\;\tilde{\rho}_{T,1}$
$\lesssim\sum_{T\in\Omega_{h}}\|\operatorname{curl}E_{\tau}\|_{\Omega_{T}}\;\tilde{\rho}_{T,1},$
@@
(26) $\tilde{\rho}(u-v_{h})\lesssim\sum_{T\in\Omega_{h}}\omega_{T,1}\;\tilde{\rho}_{T,1}+\omega_{T,2}\;\tilde{\rho}_{T,2}+\omega_{T,3}\;\tilde{\rho}_{T,3}.$
@@
(27) $\rho(u-\tilde{v}_{h})\lesssim\sum_{T\in\Omega_{h}}\tilde{\omega}_{T,1}\;\rho_{T,1}+\tilde{\omega}_{T,2}\;\rho_{T,2}+\tilde{\omega}_{T,3}\;\rho_{T,3}.$
Finally, note that the definitions of $\sigma$ and $B$ imply
$2\sigma$ $=B(u-u_{h},\tilde{u}-\tilde{u}_{h})=m(E_{\tau}-E_{\tau,h},\tilde{E}_{\tau}-\tilde{E}_{\tau,h})$
$=(\gamma(E_{\tau}-E_{\tau,h}),\tilde{E}_{\tau}-\tilde{E}_{\tau,h})_{\Omega}<1,$
where the last inequality is due to (21). Thus $21-\sigma>1/$ 2. Using this, together with (27) and (26) within the identity (23), the stated estimate follows.
@@
The next practicality involves aggregating estimators when multiple eigenvalues $\lambda_{h}^{(\ell,k)}$ are found clustered, together with their left and right eigenvectors $\tilde{u}_{h}^{(\ell,k)},u_{h}^{(\ell,k)},\;k=1,\dots,K_{\ell}$, in the $\ell$-th adaptive iteration. Then, we simply compute the element-wise maximum of the error indicators for each eigenvalue. Namely, we compute $\eta_{T}^{(\ell,k)}=\eta_{T}^{(h)}(\lambda_{h}^{(\ell,k)},u_{h}^{(\ell,k)},\tilde{u}_{h}^{(\ell,k)})$ using (28) and set
(29) $\eta_{T}^{(\ell)}=\max_{k=1,\dots,K_{\ell}}\eta_{T}^{(\ell,k)}.$
An element $T^{\prime}\in\Omega_{h}^{(\ell)}$ is marked for refinement if
(30) $\eta_{T^{\prime}}^{(\ell)}>\theta\max_{T\in\Omega_{h}}\eta_{T}^{(\ell)},$
where $\theta$ is an input refinement threshold parameter. This is used in Algorithm 1, which summarizes the adaptive strategy we implemented.
@@
Input: Initial mesh $\Omega_{h}^{(0)},$ refinement threshold $\theta$, an eigenvalue search region $D\subseteq\mathbb{C}$, and the maximum permitted number of degrees of freedom $N_{\mathrm{max}}$.
@@
4       SOLVE: Assemble the system in (18) using $\Omega_{h}^{({\ell})}$ and solve it using a sparse eigensolver to obtain a cluster of eigenvalues $\lambda_{h}^{({\ell},k)}=(Z_{h}^{({\ell},k)})^{2}$ contained in $D$, and the corresponding right and left eigenvectors $u_{h}^{({\ell},k)}$ and $\tilde{u}_{h}^{({\ell},k)}$ for $k=1,\dots,K_{\ell}$.;
5
6      ESTIMATE: Compute error indicators $\eta_{T}^{({\ell})}$ using (29) for each element $T\in\Omega_{h}^{({\ell})}$.;
7
8      MARK elements for refinement based on (30) and input $\theta$.;
9
10      REFINE marked elements in $\Omega_{h}^{({\ell})}$ (as well as surrounding elements as needed to obtain mesh conformity) to generate the next mesh $\Omega_{h}^{({\ell}+1)}$.;
@@
12      if number of degrees of freedom on $\Omega_{h}^{({\ell}+1)}>N_{\mathrm{max}}$ then
13             break;
@@
15
16return $\Omega_{h}^{({\ell})}$ and $\varLambda_{h}^{({\ell})}$;
@@
We conclude this section by specifying the parameters we used in Algorithm 1 and ensuing computations. We use NGSolve [28] for Nédélec and Lagrange finite elements and for assembly of the eigensystem. It supports curved triangles, with a specifiable order of polynomial curving, as well as high-order Nédélec and Lagrange elements. Algorithm 1 can use any sparse eigensolver. In our computations, we use the FEAST contour integral eigensolver [25, 9] (whose specific application to leaky modes is also detailed in [10, Section A.4]), which takes as input a search region enclosed by a simple closed contour (denoted by $D$ in the algorithm). We use circular contours of small enough radius centered at $\hat{\lambda}_{h}^{(n)}=\frac{1}{K_{n}}\sum_{k=1}^{K_{n}}\lambda_{h}^{(n,k)}$, i.e., these centers track changes within each adaptive iteration and $D$ may therefore vary with iteration number $\ell$. How aggressively the refinements are made is determined by the parameter $\theta$. For our computations, we used $\theta=0.75$, a value determined by trial and error to give us a good balance between localization of the refinements and a steady increase in the number of degrees of freedom (d.o.f.s). So as to represent all eigenmodes of the eigenvalue cluster on the same mesh, the algorithm uses a marking strategy based on the maximum of the error indicators generated by each eigenmode. Other marking strategies are possible, as in [33, 1, 12]. As a stopping criterion, we set the maximum number of degrees of freedom to $N_{\text{max}}=2\times 10^{6}$ in our numerical experiments.
4 Verification by Bragg fiber
@@
These prior studies make the Bragg configuration an ideal example for verifying our adaptive strategy. Because we can semi-analytically compute the exact eigenvalue $\lambda=Z^{2}$ for this geometry, we are able to study the history of actual eigenvalue errors as the adaptive mesh refinement proceeds. Moreover, starting with an unbiased mesh that disregards the above-mentioned prior knowledge of the fine-scale features in the modes, we investigate if the (deliberately blinded) adaptivity process is able to sense these features automatically and guide refinement to capture them.
@@
(b) Final mesh for the Bragg fiber after the adaptive algorithm.
@@
Figure 2: Results from the adaptive algorithm applied to a Bragg fiber.
In our numerical experiments for this fiber, we use the following settings for parameters (recalling that the parameters were defined in Section 3). The length scale in nondimensionalization is set to $L=1.5\times 10^{-5}$ m. The operating wavelength is determined by setting $k=\frac{2\pi}{1.7}\times 10^{6}\leavevmode\nobreak\ \text{m}^{-1}$. The refractive index is a piecewise-constant function whose values are $n_{\rm air}=1.00027717$ in the air regions and $n_{\rm glass}=1.43881648$ in the glass ring. The following non-dimensionalized radii are used: the air core has radius $r_{\rm core}=2.7183$; the glass ring starts at $r_{\rm core}$ and has an outer radius $r_{\rm outer}=3.385$, with a thickness of $t_{\rm ring}=0.66666667$; the outer air region starts at $r_{\rm outer}$ and extends to $r_{0}=4.385$, with a thickness of $t_{\rm air}=1.0$; and the PML starts at $r_{0}$ and extends to $r_{1}=8.05166666$, with a thickness of $t_{\rm PML}=3.66666667$. The PML strength (recall (14)) is set to $\alpha=2.0$. We set polynomial degree of $p=6$ to define the discretization (15).
@@
Next, we turn to examine the accuracy of the computed eigenvalue. The results are in Figure 3. The exact nondimensional eigenvalue we aim to approximate, obtained from semi-analytical computations, is $Z^{2}=0.80953881+0.00170153i$, and the exact eigenvalue cluster is the singleton containing it, $\varLambda=\{Z^{2}\}$. The discrete eigenvalue cluster $\varLambda_{h}$ is computed by providing a circular contour to FEAST centered around the exact eigenvalue. When the eigensolver converged, $\varLambda_{h}$ generally had two elements, matching the multiplicity of the exact eigenvalue. Let the Hausdorff distance between two sets $\Upsilon_{1}$ and $\Upsilon_{2}$ be denoted by
$d(\Upsilon_{1},\Upsilon_{2})=\max(\sup_{\gamma_{1}\in\Upsilon_{1}}\operatorname{dist}(\gamma_{1},\Upsilon_{2}),\sup_{\gamma_{2}\in\Upsilon_{2}}\operatorname{dist}(\gamma_{2},\Upsilon_{1})).$
We report the convergence of the discrete eigenvalues in $\varLambda_{h}$ to $\varLambda$ using the Hausdorff distance in Figure 3.
@@
Figure 3: Convergence history of the cluster of discrete eigenvalues to the exact eigenvalue for the Bragg fiber, during the adaptive mesh refinement.
@@
$10^{5}$$10^{6}$$10^{-7}$$10^{-5}$$10^{-3}$$10^{-1}$$10^{1}$$10^{3}$umber of degrees of freedom at the $\ell$-th iteration$|\text{CL}_{\ell}-\operatorname{CL}|$
Figure 4: Convergence history of the confinement loss (see (31)) for the Bragg fiber, during the adaptive mesh refinement.
Finally, we report on how the error estimator compares to the actual eigenvalue error (which can be computed for this example since an exact solution is available). Consider the ratio of the the global error of the computed eigenvalues to the global error estimator, i.e., at the $\ell$th iteration, let
(32) $\text{Efficiency}({\ell})=\max_{\lambda_{h}\in\varLambda_{h}^{(\ell)}}|\lambda_{h}-\lambda|\Bigg{/}\Bigg{(}\sum_{T\in\Omega_{h}^{(\ell)}}(\eta_{T}^{(\ell)})^{2}\Bigg{)}^{1/2},$
where $\lambda=Z^{2}$ is the above-mentioned known exact eigenvalue. These values, for each mesh in the history of adaptive refinements, are plotted in Figure 5. As can be seen from Figure 5, these values oscillates greatly in a preasymptotic regime, after which they stabilize somewhat, hovering well below one. On all meshes except one, these values are below one, indicating that the eigenvalue error is bounded by the computed error estimator, i.e., the estimator is quite reliable. However, we also observe that the efficiency values are not very close to the perfect value of one. Such efficiencies are comparable to previous reports [12] that lead us to not anticipate perfect efficiencies for the DWR estimator, even for textbook nonselfadjoint eigenproblems simpler than ours.
$10^{5}$$10^{5.2}$$10^{5.4}$$10^{5.6}$$10^{5.8}$$10^{6}$$10^{6.2}$$10^{-12}$$10^{-9}$$10^{-6}$$10^{-3}$$10^{0}$umber of degrees of freedom at the $\ell$-th $(\ell)$
Figure 5: Plot of ratios of eigenvalue error to the estimator (see (32)) on each mesh during the adaptive process applied to the Bragg fiber.
5 Three microstructured fibers